OpenStudy (anonymous):

WILL MEDAL AND FAN!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! A handy man knows from experience that his 29-foot ladder rests in its most stable position when the distance of its base from a wall is 1 foot farther than the height it reaches up the wall. A) How far up a wall does this ladder reach? Show your work B) How far should the base be from the wall? C) If the man needs to reach a window 25 feet high on the wall, how far from the wall should he place the base of the ladder?

OpenStudy (anonymous):

@Austin1617 @mynameisnemo @surjithayer @srizek

OpenStudy (carlyleukhardt):

x^2 + (x + 1)^2 = 29^2 x^2 + x^2 + 2x + 1 = 841 2x^2 + 2x - 840 = 0 x^2 + x - 420 = 0 x = (-1 +/- sqrt(1 + 1680)) / 2 x = (-1 +/- sqrt(1681)) / 2 x = (-1 +/- 41) / 2 x = -42/2 , 40/2 x = -21 , 20 It's a 20-21-29 foot triangle a) 20 feet b) 21 feet 29^2 - 25^2 = x^2 841 - 625 = x^2 216 = x^2 6 * sqrt(6) = x x = 14.696938456699068589183704448235 14.697 feet

OpenStudy (anonymous):

@carlyleukhardt which question does that answer?

OpenStudy (anonymous):

@carlyleukhardt ??

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

@Michele_Laino can you help with this one maybe?